4 years ago, I was half the age I will be 7 years from now. How old am I?

Answer by thejackal(72)

very easy:
Let my present age be x
such that "4 years ago" = x - 4
and "half the age I will be 7 years from now" = (x + 7)/2
thus:
x - 4 = (x + 7)/2
2x - 8 = x + 7
2x - x = 8 + 7
x = 15
proof:
4 years ago = 15 - 4 = 11
half the age I will be 7 years from now = (15 + 7)/2 = 11
QED